OpenStudy (anonymous):

when using the substitution u=root(1+x), an antiderivative of int(60x(root(1+x))dx) is

OpenStudy (dumbcow):

\[\int\limits_{}^{}60\sqrt{1+x}dx\] \[u = \sqrt{1+x} \] \[du = \frac{dx}{2\sqrt{1+x}} \rightarrow dx = 2u du\] \[=\int\limits_{}^{}120 u^{2} du\] \[=40u^{3} + C\] \[=40(1+x)\sqrt{1+x} +C\]